OpenStudy (kitkat1):

A pencil at a stationery store costs $1, and a pen costs $1.50. Stella spent $21 at the store. She bought a total of 18 items. Which system of equations can be used to find the number of pencils (x) and pens (y) she bought? (5 points) 1.5x + y = 21 x = 18y 18x + y = 21 x = 1.5y x + 1.5y = 21 x + y = 18 x + 18y = 21 x = 1.5y

OpenStudy (kitkat1):

Medals!

OpenStudy (whpalmer4):

Okay, if a pencil costs $1, and \(x\) represents the number of pencils, how do you write an expression for the amount of money spent on pencils?

OpenStudy (kitkat1):

i think like the 3rd answer

OpenStudy (whpalmer4):

well, let's figure it out. how do you write the expression for the amount spent on pencils, if a pencil costs $1 and there are \(x\) pencils?

OpenStudy (kitkat1):

Idk

OpenStudy (whpalmer4):

what is the cost of 1 pencil?

OpenStudy (kitkat1):

1$

OpenStudy (whpalmer4):

what is the cost of 2 pencils?

OpenStudy (kitkat1):

2$

OpenStudy (whpalmer4):

what is the cost of 57 pencils? (btw, the $ goes in front)

OpenStudy (whpalmer4):

you get the cost by multiplying the cost of 1 item by the number of items, right?

OpenStudy (kitkat1):

$57

OpenStudy (kitkat1):

Im pretty sure the answer is the 3rd one

OpenStudy (whpalmer4):

so the cost of \(x\) pencils when pencils cost $1 each is \(\$1*x\) or \(1x\) or \(x\) if we assume the price is in dollars. Pens cost $1.50 each. If you have \(y\) pens, how much is that?

OpenStudy (kitkat1):

1.5y

OpenStudy (whpalmer4):

I agree the answer is C: I'm trying to solidify your understanding of this so you never have to ask again, you'll just know the right answer. Worth a minute or two of your time, right? Yes, 1.5y is the expression for the cost of the pens. We know that we spent $21 on \(x\) pencils and \(y\) pens, so how do we write that in an equation?

OpenStudy (kitkat1):

x + 1.5y=21

OpenStudy (whpalmer4):

Very good. And if we know that we have a total of 18 items, how do we write that?

OpenStudy (kitkat1):

x + y = 18

OpenStudy (kitkat1):

Can u help me with one more?

OpenStudy (whpalmer4):

Excellent. For a medal of your own, how many pens and pencils?

OpenStudy (kitkat1):

18 and 21

OpenStudy (kitkat1):

The equation 2.2x – 4.4 = 1.8x can be transformed to form which of the following expressions? 4.0x = 4.4 0.4x = –4.4 0.4x = 4.4 4.0x = –4.4 Can u help me with this one?

OpenStudy (whpalmer4):

No, 18 pens and 21 pencils doesn't add up to 18 items, does it?

OpenStudy (kitkat1):

31

OpenStudy (whpalmer4):

Come on, you aren't thinking. The total number of pens + pencils is 18. The total cost is $21. Well, whatever. Let's do your other problem. \[2.2x – 4.4 = 1.8x\] What happens if you subtract \(1.8x\) from both sides?

OpenStudy (kitkat1):

2.2x - 2.6

OpenStudy (kitkat1):

Now what?

OpenStudy (kitkat1):

so 0.4x in the beginning

OpenStudy (whpalmer4):

\[2.2x-4.4-1.8x=1.8x-1.8x\]\[2.2x-1.8x-4.4 = 0x\]\[0.4x-4.4=0\] Now we add 4.4 to both sides: \[0.4x-4.4+4.4=0+4.4\]\[0.4x=4.4\]

OpenStudy (kitkat1):

Thank you!